Durable press cotton fabrics are treated to improve their recovery from wrinkles after washing. Wrinkle recovery angle me
Answer #1

The given data as, permafresh 15 Hylite 16 18 13 16 jo 19 20 12 1 Hese, Foo Permafrest, ni-6 X 7 1 2 12.833 Si= hi (zxj2 nă?)Summamy, permafresh ni = 6 X = 12.883 Si = 2.3166 5,² - 5.3668 Hylite 12=6 72- 17.933 S2 = 2.1602 S2 = 4.66 66 TO find, gol c